OpenStudy (anonymous):

Which of the following describes the correct process for solving the equation 3x + 6 = 21 and arrives at the correct solution?

OpenStudy (anonymous):

1) subtract 6 from both sides 2) divide both sides by 3

OpenStudy (anonymous):

3x+6=21 1) subtract 6 from both sides 3x+6-6=21-6 3x=15 2) divide both sides by 3 3x/3= 15/3 x=5